# Program evaluation with remotely sensed outcomes

### v2

@@ 4.3 Robustness to mis-specification
Assumption 4.1 does not require any complexity restriction, nor any rate of convergence. The moment restriction in Theorem 3.1 is infinite-order Neyman orthogonal (Mackey et al.,, 2018; Chen et al.,, 2020), so Algorithm 4.1 enjoys the best of both worlds: no complexity restriction (Chernozhukov et al.,, 2018, 2023), and no rate requirement (Chamberlain,, 1987; Newey,, 1993).
@@
Proposition 4.1 (Inference with known counts).
Suppose Theorem 3.1’s conditions and Assumption 4.1 hold. If the marginal probabilities are known and bounded away from zero, then $n^{1/2}(\hat{\theta}-\theta)\rightsquigarrow\mathcal{N}\left(0,\frac{\mathbb{E}\{(\Delta^{e}-\theta\Delta^{o})^{2}\widetilde{H}(R)^{2}\}}{[\mathbb{E}\{\Delta^{o}\widetilde{H}(R)\}]^{2}}\right)$. Moreover, if $\widetilde{H}(R)=H^{*}(R)$, then $\hat{\theta}$ is semiparametrically efficient for $\theta$ satisfying the conditional moment in Theorem 3.1.
Proposition 4.2 (Inference with unknown counts).
Suppose Theorem 3.1’s conditions and Assumption 4.1 hold. If the marginal probabilities and their counting estimators are bounded away from zero, then $n^{1/2}(\hat{\theta}-\theta)\rightsquigarrow\mathcal{N}\left(0,\frac{V}{[\mathbb{E}\{\Delta^{o}\widetilde{H}(R)\}]^{2}}\right)$, where $V$ is defined in Lemma D.6.
When marginal probabilities are known, the asymptotic variance is standard (Proposition 4.1). When marginal probabilities are unknown, the asymptotic variance is weighted by them (Proposition 4.2). For the latter, it is simpler to use a bootstrap than an analytic variance estimator.
Theorem 3.2 allows direct effects of the treatment on the remotely sensed variable, as long as treatment is non-missing and non-deterministic in the observational sample. We have already derived the conditional moment equations. Estimation and inference remain essentially the same.
@@ 5 Program evaluation using satellite images
Across the exercises, we use data from an experiment analyzed in Muralidharan et al., (2016, 2023), illustrated in Figure 1. The authors collect data in an experimental sample and an observational sample of villages in Andra Pradesh, India. The treatment $D$ is the early introduction of Smartcards, which are a biometrically authenticated payments infrastructure. The outcome $Y$ is a measure of village-level poverty. See Appendix G for details on the experiment.
We use the geographic coordinates of each village to extract a remotely sensed variable $R$, from free, open-access databases. The RSV concatenates luminosity (a scalar) (Asher et al.,, 2021) and satellite images (a high-dimensional, pre-trained embedding vector in $\mathbb{R}^{4000}$) (Rolf et al.,, 2021). These variables have been extensively validated as predictors of poverty (Henderson et al.,, 2011; Jean et al.,, 2016; Stoeffler et al.,, 2016; Michaels et al.,, 2021; Huang et al.,, 2021; Sherman et al.,, 2023). Our framework allows us to test their relevance for program evaluation, similar to a “first stage” exercise in instrumental variable analysis. In our semi-synthetic exercises, we will evaluate how well we can conduct program evaluation using this real RSV.
In terms of our identification framework, the semi-synthetic settings have incomplete cases (Assumption 2.3(ii)): the treatment is missing in the observational sample. The settings also have some experimental outcomes (Remark 3.3). We will be explicit about sample definitions below.
@@ 5.1 Our method outperforms current practice across effect sizes and sample sizes
We impose all of our assumptions in a calibrated, synthetic data generating process (DGP). First, we simulate the binary treatment $D$ as a fair coin toss (Assumption 2.1). If $D=0$, we simulate the binary outcome $Y$ as a weighted coin toss, calibrated to the empirical probability of $Y=1$ among untreated experimental units in the real data. If $D=1$, we simulate $Y$ as a different weighted coin toss, calibrated to the empirical probability of $Y=1$ among treated experimental units in the real data.
In this baseline version of the DGP, the synthetic treatment effect is calibrated to the real one: $0.07$. We consider additional synthetic treatment effect values $\theta=0.07+\tau$ by augmenting the probability of $Y=1$ when $D=1$ for alternative values of $\tau$.
Next, we draw RSV values from the real data. If $Y=0$, we draw $R$ from the empirical distribution of $R|Y=0$ in the real experimental data. Likewise, we draw $R$ when $Y=1$. This imposes no direct effects (Assumption 2.3(ii)). For computational feasibility, we use luminosity and only the initial $1,000$ satellite image features as our RSV.
[…]
@@
Figures 4 and 5 demonstrate that our method outperforms common practice across treatment effect values and sample sizes. We consider treatment effect values $\theta=0.07+\tau$ with $\tau\in\{0,0.1,...,0.5\}$ and sample sizes $n\in\{1000,2000,3000\}$. By varying these aspects of the synthetic DGP, we evaluate relative performance for different signal-to-noise ratios.
We compare the bias and root mean square error of the two methods. Whereas the bias of our method is always small and vanishing with sample size, the bias of the common practice is typically very large and constant across sample sizes. Common practice has positive or negative bias for the treatment effect. While the variance of our method is similar to the variance of the common practice, our method’s large improvement in bias translates into similar improvement in mean square error when the sample size is large enough.
This exercise has clear consequences for empirical practice. If the goal is to conduct valid inference on the treatment effect, common practice can be misleading due to large bias that does not vanish as the sample size increases. It can return the wrong sign, as foreshadowed by Proposition 3.1. By contrast, our method reports unbiased estimates, with valid asymptotic inference.
@@ 5.2 Our method recovers the true effect with randomly missing outcomes
While our first exercise involves synthetic treatment effects, our second exercise involves real treatment effects. We now ask: compared to an unbiased benchmark, how does our method perform? The benchmark is the difference-in-means estimate and confidence interval an economist would obtain if they could observe both treatments and outcomes in the experiment. Our method gives an estimate and confidence interval using treatments and RSVs in an experiment, and outcomes and RSVs in an observational sample. [^11] Following Remark 3.3, we also observe some outcomes in the experiment, as described below.
We conduct this exercise with three possible poverty measurements: (i) is the village in the bottom quartile of villages for per capita consumption, (ii) does a village have only low income households, and (iii) does a village have only low and middle income households. While (i) is a measure of average consumption, (ii) and (iii) describe the income distribution using formal categories from Indian administrative data; see Appendix G for details. [^12] We used the low consumption outcome (i) in the first semi-synthetic exercise.
In this exercise, we now use the full RSV: luminosity, and the full, 4000-dimensional embedding of satellite images.
In a semi-synthetic way, we classify villages into the samples described in Assumption 2.3(ii) and Remark 3.3. We classify the real observational villages as $\tilde{S}=o$, for which we observe $(Y,R)$. Among the real experimental villages, we randomly classify some as $\tilde{S}=e$, for which we observe $(D,R)$, and some as $\tilde{S}=\{e,o\}$, for which we observe $(D,Y,R)$. In other words, for real experimental villages, we randomly delete half of their outcomes.
This exercise satisfies the key assumptions of our framework. The real treatment variable was randomized in the real experiment (Assumption 2.1). Stability of the RSV conditional distribution is plausible, as demonstrated in Figure 2 (Assumption 2.2). Finally, because the real data have incomplete cases, we must argue that the treatment only affects the RSV via the outcome (Assumption 2.3(ii)). As supporting evidence, Figures 11 and 13 show that the distribution of $R|Y=y,D=1$ visually matches the distribution of $R|Y=y,D=0$.
Another requirement in our framework is that the RSV is relevant: $\mathbb{E}\{H(R)\Delta^{o}\}\neq 0$ in Corollary 3.1. Intuitively, the representation of the RSV $H(R)$ used for inference should be correlated with outcome variation $\Delta^{o}$. This requirement is testable; for our learned representation in Algorithm 4.1, we can test whether the empirical analogue $\mathbb{E}_{n}\{\widehat{H}(R)\widehat{\Delta}^{o}\}$ is significantly nonzero. Figure 6 confirms that our RSV is relevant.
@@
We further interpret and compare our learned representation of the satellite image in Appendix G. Our optimal representation is based on three predictions: the outcome, treatment, and sample indicator given the RSV. By contrast, common practice only uses the prediction of the outcome given the RSV. Interestingly, our optimal representation allows for extrapolation, i.e. negative weights on villages, whereas common practice does not.
@@
Figure 7 (“synthetic samples”) shows that our method recovers the treatment effect estimated by the unbiased benchmark in this exercise. For each poverty outcome, our treatment effects have the same signs and magnitudes as the benchmark, i.e. as if the economist could observe treatments and outcomes in the experiment. These effects are consistent with the findings in Muralidharan et al., (2023): early adoption of Smartcards reduces poverty. Compared to the benchmark, our confidence intervals are similar, and sometimes shorter; efficiently using the additional information in RSVs can boost statistical power. [^13] Assumption 2.3(ii) is an additional restriction that may improve asymptotic precision. Relative magnitudes of standard errors may also reflect finite sample estimation error.
These findings have a practical implication: our method may allow economists to significantly reduce survey costs. In this exercise, we mimic what would happen if an economist paid surveyors to collect outcomes for half of the experimental villages, and relied on free satellite images for the other half. We can calculate the savings, compared to paying surveyors to collect outcomes for all of the experimental villages. If surveying each individual in a village conservatively costs \$$\$0.50$, our results suggest savings of \$$\$3.6$ million. [^14] Table 5 summarizes the number of experimental villages and the populations in experimental villages, which we use to calculate savings. This calculation is extremely conservative; Viviano and Rudder, (2024) find that phone surveys in Pakistan cost \$$\$7$ per individual, rather than \$$\$0.50.$
@@ 5.3 Our method recovers the true effect in a realistic setting
Finally, as our third exercise, we use the data of Muralidharan et al., (2016, 2023) as realistically as possible. The third exercise is identical to the second exercise, except that we classify villages in a more realistic way. As before, we classify the real observational villages as $\tilde{S}=o$, for which we observe $(Y,R)$. Among the real experimental villages, we now classify the treated ones as $\tilde{S}=e$, for which we observe $(D,R)$, and the untreated ones as $\tilde{S}=\{e,o\}$, for which we observe $(D,Y,R)$. In other words, for real experimental villages, we systematically delete the outcomes of all treated villages. Methods based on missingness-at-random provide no guidance in this setting.
As before, this exercise appears to satisfy our identifying assumptions. The real treatment variable was randomized in the real experiment (Assumption 2.1). The conditional distribution of the RSV appears stable in Figure 2 (Assumption 2.2). There do not appear to be direct effects in Figures 11 and 13 (Assumption 2.3(ii)). Moreover, the RSV is relevant in Figure 6 (Corollary 3.1).
In this more challenging exercise, we find qualitatively similar results. Our learned representations allow efficient extrapolation in Figure 14. Our method recovers similar estimates and confidence intervals to the benchmark in Figure 7 (“real samples”), consistent with prior work.

### v3

@@ 4.3 Robustness to Misspecification
Assumption 4 does not require any complexity restriction, nor any rate of convergence. The moment restriction in (2) is infinite-order Neyman orthogonal (Mackey et al., 2018; Chen et al., 2020), so Algorithm 1 enjoys the best of both worlds: no complexity restriction (Chernozhukov et al., 2018, 2023), and no rate requirement (Chamberlain, 1987; Newey, 1993). Even if all three of its predictions are misspecified, Algorithm 1 still delivers valid $n^{-1/2}$ inference, which is a strong form of robustness to misspecification.
@@
Proposition 2 (Inference with known counts).
Suppose Theorem 1’s conditions and Assumption 4 hold. If the marginal probabilities are known and bounded away from zero, then $n^{1/2}(\hat{\theta}-\theta)\rightsquigarrow\mathcal{N}\left(0,\frac{\mathbb{E}\{(\Delta^{e}-\theta\Delta^{o})^{2}\widetilde{H}(R)^{2}\}}{[\mathbb{E}\{\Delta^{o}\widetilde{H}(R)\}]^{2}}\right)$. Moreover, if $\widetilde{H}(R)=H^{*}(R)$, then $\hat{\theta}$ is semiparametrically efficient for $\theta$ satisfying (2) with marginal probabilities known to the researcher.
Proposition 3 (Inference with unknown counts).
Suppose Theorem 1’s conditions and Assumption 4 hold. If the marginal probabilities and their counting estimators are bounded away from zero, then $n^{1/2}(\hat{\theta}-\theta)\rightsquigarrow\mathcal{N}\left(0,\frac{V}{[\mathbb{E}\{\Delta^{o}\widetilde{H}(R)\}]^{2}}\right)$, where $V$ is defined in Lemma D.6.
When marginal probabilities are known, the asymptotic variance is standard (Proposition 2). In this case, (2) matches the conditional moment of Chamberlain (1987), so the semiparametric efficiency bound is known.
When marginal probabilities are unknown, the asymptotic variance is weighted by them (Proposition 3). The asymptotic variance in Proposition 3 differs from the one in Proposition 2 because estimation of the marginal probabilities introduces an additional estimation error of order $n^{-1/2}$ that we must take into account for asymptotic inference. In this case, it is simpler to use a bootstrap than an analytic variance estimator.
Theorem 2 allows direct effects of the treatment on the remotely sensed variable, as long as treatment is non-missing and non-deterministic in the observational sample. We have already derived the conditional moment equations. Estimation and inference are similar.
Remark 6 (Inference with covariates).
For clarity of exposition, this section has focused on the setting without covariates $X$.
When $X$ is discrete with a finite support, estimation and inference are straightforward: apply Algorithm 1 within each covariate stratum, then average over $X$. See Appendix E.3.
When $X$ is continuous with a low dimension, there are two main modifications: (i) estimate the conditional probabilities $\Pr(D=d,S=e\mid X=x)$ and $\Pr(Y=y,S=o\mid X=x)$ as smooth functions of $x$; (ii) estimate the conditional average treatment effect $\theta(x)$ via local regression or series methods, before averaging over $X$. Appendix E.4 discusses this setting.
Crucially, the main robustness property—that $n^{-1/2}$ inference requires no rate condition on the learned representation $\widehat{H}(x,R)$—continues to hold in both settings. This justifies the use of complex machine learning algorithms for RSV representations.
@@ 5 Program Evaluation using Satellite Images
Across the exercises, we use data from an experiment analyzed in Muralidharan et al. (2016, 2023), illustrated in Figure 1. The authors collect data in an experimental sample and an observational sample of villages in Andhra Pradesh, India. The treatment $D$ is the early introduction of Smartcards, which are a biometrically authenticated payments infrastructure. The outcome $Y$ is a measure of village-level poverty in administrative data from 2012 to 2013 (after the experiment was deployed). See Appendix H for details on the experiment.
We use the geographic coordinates of each village to extract a remotely sensed variable $R$, from free, open-access databases. The RSV concatenates nighttime luminosity measures from 2012 to 2020 (a vector in $\mathbb{R}^{50}$) (Asher et al., 2021) and satellite images from 2019 (a high-dimensional, pre-trained embedding vector in $\mathbb{R}^{4000}$) (Rolf et al., 2021). These variables have been extensively validated as predictors of poverty (Henderson et al., 2011; Jean et al., 2016; Stoeffler et al., 2016; Michaels et al., 2021; Huang et al., 2021; Sherman et al., 2023). Our framework allows us to test their relevance for program evaluation, similar to a “first stage” exercise in instrumental variable analysis. In our semi-synthetic exercises, we will evaluate how well we can conduct program evaluation using this real RSV.
In terms of our identification framework, the semi-synthetic settings have incomplete cases (Assumption 3(ii)): the treatment is missing in the observational sample. The settings also have some experimental outcomes (Remark 4). We will be explicit about sample definitions below. Appendix H contains a complete description of the data.
@@ 5.1 Our Method Outperforms Current Practice across Effect Sizes and Sample Sizes
We impose all of our assumptions in a calibrated, synthetic data-generating process (DGP). First, we simulate the binary treatment $D$ as a fair coin toss (Assumption 1). If $D=0$, we simulate the binary outcome $Y$ as a weighted coin toss, calibrated to the empirical probability of $Y=1$ among untreated experimental units in the real data. If $D=1$, we simulate $Y$ as a different weighted coin toss, calibrated to the empirical probability of $Y=1$ among treated experimental units in the real data.
In this baseline version of the DGP, the synthetic treatment effect is calibrated to the real one: 0.07-0.07. We consider additional synthetic treatment effect values $\theta=-0.07+\tau$ by augmenting the probability of $Y=1$ when $D=1$ for alternative values of $\tau$.
Next, we draw RSV values from the real data. If $Y=0$, we draw $R$ from the empirical distribution of $R|Y=0$ in the real experimental data. Likewise, we draw $R$ when $Y=1$. This imposes no direct effects (Assumption 3(ii)). For computational feasibility in this exercise, we use luminosity and only the initial $1,000$ satellite image features as our RSV.
Finally, to mimic a setting with missing outcomes, we delete $Y$ if $D=1$. In terms of Remark 4, we set $\tilde{S}=e$ when $D=1$ and $\tilde{S}=\{e,o\}$ when D=0D=0.131313For simplicity, we do not have $\tilde{S}=o$. This could be easily done: if $D=0$, toss a coin to determine whether $\tilde{S}=\{e,o\}$ or $\tilde{S}=o$. With or without $\tilde{S}=o$, a direct extension of Theorem 1 applies: within the identifying formula given there, simply replace $S=e$ with $e\in\tilde{S}$ in $\Delta^{e}$, and similarly replace $S=o$ with $o\in\tilde{S}$ in $\Delta^{o}$. This imposes stability across samples (Assumption 2). Recent methods that use machine learning predictions as outcomes, e.g., prediction-powered inference (Angelopoulos et al., 2023), provide no guidance for this DGP, because no treated unit has an observed outcome.
@@
Figures 4 and 5 demonstrate that our method outperforms common practice across treatment effect values and sample sizes. We consider treatment effect values $\theta=-0.07+\tau$ with $\tau\in\{0,0.1,\ldots,0.5\}$ and sample sizes $n\in\{1000,2000,3000\}$. By varying these aspects of the synthetic DGP, we evaluate relative performance for different signal-to-noise ratios.
We compare the bias and root mean square error of the two methods. Whereas the bias of our method is always small and vanishing with sample size, the bias of the common practice is typically very large and constant across sample sizes. Common practice has positive or negative bias for the treatment effect, consistent with Proposition 1. While the variance of our method is similar to the variance of the common practice, our method’s large improvement in bias translates into similar improvement in mean square error when the sample size is large enough.
This exercise has clear consequences for empirical practice. If the goal is to conduct valid inference on the treatment effect, common practice can be misleading due to large bias that does not vanish as the sample size increases. It can return the wrong sign, as characterized by Proposition 1. By contrast, our method reports unbiased estimates, with valid asymptotic inference.
@@ 5.2 Our Method Recovers the True Effect with Randomly Missing Outcomes
While our first exercise involves synthetic treatment effects, our second exercise involves real treatment effects. We now ask: compared to an unbiased benchmark, how does our method perform? The benchmark is the difference-in-means estimate and confidence interval an economist would obtain if they could observe all treatments and outcomes in the experiment. Our method gives an estimate and confidence interval using treatments and RSVs in an experiment, and outcomes and RSVs in an observational sample. [^14] Following Remark 4, we also observe some outcomes in the experiment, as described below.
We conduct this exercise with three possible poverty measurements: (i) is the village in the bottom quartile of villages for per capita consumption, (ii) does a village have only low income households, and (iii) does a village have only low and middle income households. While (i) is a measure of average consumption, (ii) and (iii) describe the income distribution using formal categories from Indian administrative data; see Appendix H for details. [^15] We used the low consumption outcome (i) in the first semi-synthetic exercise.
In this exercise, we now use the full RSV: the $50$ luminosity measures, and the full $4000$-dimensional embedding of satellite images.
In a semi-synthetic way, we classify villages into the samples described in Assumption 3(ii) and Remark 4. We classify the real observational villages as $\tilde{S}=o$, for which we observe $(Y,R)$. Among the real experimental villages, we randomly classify some as $\tilde{S}=e$, for which we observe $(D,R)$, and some as $\tilde{S}=\{e,o\}$, for which we observe $(D,Y,R)$. In other words, for real experimental villages, we randomly delete half of their outcomes.
This exercise satisfies the key assumptions of our framework. The real treatment variable was randomized in the real experiment (Assumption 1). Stability of the RSV conditional distribution is plausible, as demonstrated in Figure 2 (Assumption 2). Finally, because the real data have incomplete cases, we must argue that the treatment only affects the RSV via the outcome (Assumption 3(ii)). As supporting evidence, Figures H.1 and H.3 show that the distribution of $R|Y=y,D=1$ visually matches the distribution of $R|Y=y,D=0$.
Another requirement in our framework is that the RSV is relevant: $\mathbb{E}\{H(R)\Delta^{o}\}\neq 0$ in Corollary 1. Intuitively, the representation of the RSV $H(R)$ used for inference should be correlated with outcome variation $\Delta^{o}$. This requirement is testable; for our learned representation in Algorithm 1, we can test whether the empirical analogue $\mathbb{E}_{n}\{\widehat{H}(R)\widehat{\Delta}^{o}\}$ is significantly nonzero. Figure 6 confirms that our RSV is relevant.
@@
We further interpret and compare our learned representation of the satellite image in Appendix H. Our optimal representation is based on three predictions: the outcome, treatment, and sample indicator given the RSV. By contrast, common practice only uses the prediction of the outcome given the RSV. Interestingly, our optimal representation allows for extrapolation, i.e. negative weights on villages, whereas common practice does not.
@@
Figure 7 (“synthetic samples”) shows that our method recovers the treatment effect estimated by the unbiased benchmark in this exercise. For each poverty outcome, our treatment effects have the same signs and similar magnitudes as the benchmark, i.e. as if the economist could observe all treatments and outcomes in the experiment. These effects are consistent in sign with the findings in Muralidharan et al. (2023): early adoption of Smartcards reduces poverty. Compared to the benchmark, our confidence intervals are similar, and sometimes shorter; efficiently using the additional information in RSVs can boost statistical power. [^16] Assumption 3(ii) is an additional restriction that may improve asymptotic precision. Relative magnitudes of standard errors may also reflect finite sample estimation error.
These findings have a practical implication: our method may allow economists to significantly reduce survey costs. In this exercise, we mimic what would happen if an economist paid surveyors to collect outcomes for half of the experimental villages, and relied on free satellite images for the other half. We can calculate the savings, compared to paying surveyors to collect outcomes for all of the experimental villages. If surveying each individual in a village conservatively costs $\mathdollar 0.50$, our results suggest savings of $\mathdollar 3.6$ million. [^17] Table H.1 summarizes the number of experimental villages and the populations in experimental villages, which we use to calculate savings. This calculation is highly conservative; Viviano and Rudder (2024) find that phone surveys in Pakistan cost $\mathdollar 7$ per individual, rather than $\mathdollar 0.50.$
@@ 5.3 Our Method Recovers the True Effect in a Realistic Setting
The third exercise is identical to the second exercise, except that we classify villages in a more realistic way. As before, we classify the real observational villages as $\tilde{S}=o$, for which we observe $(Y,R)$. Among the real experimental villages, we now classify the treated ones as $\tilde{S}=e$, for which we observe $(D,R)$, and the untreated ones as $\tilde{S}=\{e,o\}$, for which we observe $(D,Y,R)$. In other words, for real experimental villages, we systematically delete the outcomes of all treated villages. Methods based on missingness-at-random provide no guidance in this setting.
As before, this exercise appears to satisfy our identifying assumptions. The real treatment variable was randomized in the real experiment (Assumption 1). The conditional distribution of the RSV appears stable in Figure 2 (Assumption 2). There do not appear to be direct effects in Figures H.1 and H.3 (Assumption 3(ii)). The RSV is relevant in Figure 6 (Corollary 1).
In this more challenging exercise, we find qualitatively similar results. Our method recovers comparable estimates and confidence intervals to the benchmark in Figure 7 (“real samples”). Our learned representations allow efficient extrapolation in Figure H.4.
We conclude this discussion with two final remarks on the interpretation of our semi-synthetic exercises.
Remark 7 (Spillovers).
In the first exercise, we compare our method to the true value of the treatment effect in Definition 1. There are no spillovers by design.
In the second and third exercises, we compare our method to a natural benchmark: a difference-in-means estimate an economist would obtain if they could observe all treatments and outcomes in the experiment. If there are no spillovers, then we are comparing our method to an unbiased estimate of the treatment effect in Definition 1. If there are spillovers, then we are comparing to an unbiased estimate of a reasonable estimand: a difference-in-means of expected outcomes, indexing by each unit’s treatment status and averaging over other all other units.
In Appendix I, we consider spillovers more carefully, by introducing additional notation. We discuss how our method can be adapted to estimate the global average treatment effect, a common parameter of interest in models with spillovers.
Remark 8 (Timing).
Outcomes were measured at the end of the experiment in 2013, so we study the effect on poverty in 2013. Meanwhile, RSVs were measured through 2020. Appendix I confirms that it is valid to use RSVs measured after the outcomes. The auxiliary sample links outcomes from 2013 with RSVs through 2020, so the estimand and estimator remain unchanged despite these differences in timing.